### v3

@@ 6.3 Effects of ETS Implementation
CV per capita (\$) 37.09 54.96 13.06 56.93 69.33 8.15
CV (%) 2.374 0.590 0.857 0.865 0.201 0.348
@@ 7 Sensitivity Analysis
We perform sensitivity analyses to evaluate alternative emission cap scenarios and to examine the influence of key parameters—climate damages and emission abatement—on our simulation results.
7.1 Alternative Emission Cap Scenarios
Along with the baseline emission cap scenario, we analyze simulation results for the noncooperative model with alternative emission cap paths, defined by the net zero emission targets in 2050, 2070, and 2090, where all regions achieve net zero emissions by the specified year. The top-left and top-right panels in Figure 6 display the emission permit prices and expected temperature increases under different emission cap scenarios. Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,539 per ton of carbon in 2049, and the temperature rise is restricted to 1.62°C by the end of this century. Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$514 and \$335 per ton of carbon in 2049, respectively. In these scenarios, the temperature rise by the end of the century is expected to reach 1.80°C and 1.98°C above the pre-industrial level, respectively. This result shows that the global target of restricting the temperature rise to 1.5°C is unattainable in a noncooperative world, even under the most restrictive net zero 2050 scenario, suggesting that stronger measures are needed to effectively regulate global emissions.
The bottom-left and bottom-right panels in Figure 6 show the regional MAC and SCC, taking the US as an example. The comparison of the regional SCC and MAC for all other regions are available in Appendix A.4.3, which show the same patterns as the US. Our results show that stricter emission caps lead to higher MACs. Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,471 per ton of carbon in 2048, compared to the peak of \$1,322 in 2086 under the net zero 2090 scenario. This is because more rigorous emission caps imposed on each region entail additional abatement efforts, resulting in a higher MAC. We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$226 per ton of carbon in 2050 in the net zero 2050 scenario, while it is \$270 in 2050 in the net zero 2090 scenario. This is because the permit price in the net zero 2050 scenario grows more quickly over time and even faster than the MAC (recall that the optimal carbon tax is equal to the gap between the MAC and permit price when emissions are strictly positive). Our finding suggests that under a stricter regulation on emission caps, carbon taxation becomes a less important policy instrument to control emissions.
7.2 Alternative Climate Damages and Abatement Costs
@@
Lastly, recognizing that climate damages and emission abatement costs are key drivers of our model outcomes, including permit prices, MAC, and the SCC, we conduct sensitivity analyses on the parameters for climate damages ($\pi_{1,i}$, $\pi_{2,i}$) and emission abatement ($b_{1,i}$, $b_{2,i}$, $b_{3,i},$ $b_{4,i}$). While we incorporate climate damage projections from Burke et al., (2018) as our baseline model simulation, we additionally consider projections from Kahn et al., (2021) and Nordhaus, 2010a . Specifically, we calibrate the climate damage parameters to match the projections from Kahn et al., (2021) and directly adopt the parameters from Nordhaus, 2010a . Figure 7 demonstrates the key economic and climate outcomes under different climate damage assumptions. The baseline climate damage parameters from Burke et al., (2018) result in slightly lower emission permit prices, reaching \$2,059 by 2069, compared to \$2,282 under the damage parameters estimated from Kahn et al., (2021) and \$2,294 under the damage parameters from Nordhaus, 2010a . With the same emission cap constraints imposed, the global temperature outcomes remain unaffected despite variations in the damage parameters. For the US, the baseline climate damage estimation from Burke et al., (2018) leads to higher MAC and higher SCC, indicating that the baseline marginal climate damages are relatively higher than those projected by Kahn et al., (2021) and Nordhaus, 2010a . However, regional heterogeneity exists; for example, Russia experiences negative SCC under the baseline parameter values, while experiencing small but positive SCC under the parameter values estimated from Kahn et al., (2021) and Nordhaus, 2010a . See Appendix A.4.4 for a comparison across all 12 regions.
@@
For our sensitivity analysis of the emission abatement cost parameters, we consider the parameter values from Nordhaus, 2010a in addition to the baseline parameter values calibrated from Ueckerdt et al., (2019), both of which share the same functional form of abatement cost. As shown in Figure 8, the permit price rises to \$1,767 by 2069 under the abatement cost estimate of Nordhaus, 2010a , approximately 85% of the permit price projected under the baseline scenario. Despite the lower permit price, the global temperature increase remains similar to the baseline simulation, reaching 1.70°C by the end of the century, due to the emission cap constraints. The lower emission permit price under the abatement cost estimate of Nordhaus, 2010a reflects lower MAC, as illustrated with the US case in Figure 8, with similar patterns observed across most regions (see Appendix A.4.5). Lastly, the SCC is also lower under Nordhaus, 2010a parameters, with the SCC of the US reaching \$101 by 2100—just 32% of the baseline scenario.
@@ 8 Conclusion
In this work, we build a dynamic multi-region model of climate and the economy with a global emission cap-and-trade system. In our integrated assessment framework, 12 aggregated regions are allocated emission caps in line with the emission targets of the NDCs and net zero commitments, as established under the Paris Agreement and the Glasgow Pact. We solve for the market prices of emission permits under the dynamic Nash equilibrium in a noncooperative setting. The permit prices are endogenously determined by demand and supply of emission permits in the global permits market, reflecting regional heterogeneity in future productivity growth, abatement technologies, climate damage, and population growth. We further show that carbon taxation needs to be introduced under the global ETS as a complementary policy to help the social planner of each region achieve their optimal emissions. For strictly positive emissions, we show both theoretically and numerically that the optimal carbon tax is equal to the difference between the regional MAC and the market price of permits.
This work has several policy implications. First, our results indicate that the current global target of restricting the global temperature rise to 1.5°C above pre-industrial level by 2100 is unattainable under noncooperation, with the current emission commitments outlined in the Paris Agreement and the Glasgow Pact. Our findings suggest that more stringent emission reduction targets and global cooperation are needed to curb the trend of rising global temperature. Second, our baseline simulation shows that the current emission commitments lead to excess emission permit supply in the initial years, resulting in permit prices of zero. This finding suggests that effective implementation of the global ETS requires stricter emission caps so that the global supply of permits does not exceed the global demand of permits. Third, we demonstrate that the global ETS is not a pure substitute for the carbon tax, but instead that both can be implemented together to improve the efficiency of global climate policies.
@@ A.4.2 Model Comparison of ETS Implementation
Figure A.8 compares regional emissions between two cases under noncooperation with the baseline emission caps: (i) with the ETS, (ii) without the ETS.
Figure A.8: Comparison of the ETS implementation: regional emissions.
@@
Figure A.9: Comparison of the ETS implementation: regional MAC.
@@
Figure A.10: Comparison of the ETS implementation: regional SCC.
A.4.3 Sensitivity Analysis over Emission Caps
@@
Figure A.12: Comparison of regional SCC under different.
A.4.4 Sensitivity Analysis over Climate Damage Parameters
In Figure A.13, we compare the MAC for the noncooperative model with the ETS under alternative values of the climate damage parameters ($\pi_{1,i}$, $\pi_{2,i}$), based on climate damage projections from Kahn et al., (2021) and Nordhaus, 2010a alongside the baseline climate damage projection from Burke et al., (2018).
@@ A.4.4 Sensitivity Analysis over Climate Damage Parameters
A.4.5 Sensitivity Analysis over Abatement Cost Parameters
@@
A.4.6 Sensitivity over TFP Growth Rates

### v4

@@ 7.3 Effects of ETS Implementation
CV per capita (\$) 22.94 -36.03 -5.40 66.68 109.69 -18.93
CV per capita (%) 1.469 -0.387 -0.354 1.001 0.317 -0.806
To isolate the impact of the ETS, we adjust the emission cap $\overline{E}_{i,t}$ in the noncooperative model with the ETS to $\overline{E}^{\prime}_{i,t}$, which is the optimal level of net emissions obtained from the noncooperative model with the emission caps $\overline{E}_{i,t}$ but without the ETS. We then compare economies with and without the ETS under the emission cap $\overline{E}^{\prime}_{i,t},$ ensuring two economies have the same pathways of global net emissions and temperature. Table 2 shows that the welfare impacts of the global ETS are strictly positive for all regions, irrespective of whether they are permit sellers or buyers, with considerable heterogeneity in the magnitude of these effects across the regions. For example, in the United States, the CV per capita is \$116.18, which is a relatively small fraction of per capita consumption (0.267%). Among all regions, Russia still experiences the largest welfare improvement, with a CV per capita of \$235.17, equivalent to 2.850% of its per capita consumption, though its welfare gains are smaller than in the pre-adjustment analysis. This is because Russia, a country with a negative SCC, benefits from higher temperatures, meaning that its welfare gains in the pre-adjustment analysis reflect both climate-induced benefits and ETS implementation. Conversely, Africa and non-OECD Asia experience the smallest welfare gains from the global ETS, indicating that additional climate damages from higher temperatures contributed to the negative welfare effects observed in the pre-adjustment analysis. Overall, the results in the post-adjustment analysis highlight that the principle of gains from trade applies to emissions trading as well, driven by efficiency gains achieved through the reallocation of emission abatement efforts across regions.
Table 2: Welfare effects of the ETS implementation: post-emission cap adjustment.
US EU Japan Russia Eurasia China
CV per capita (\$) 116.18 100.26 38.60 235.17 57.49 34.62
CV per capita (%) 0.267 0.399 0.136 2.850 1.148 0.346
India MidEast Africa LatAm OHI OthAs
CV per capita (\$) 33.34 42.73 16.82 75.30 100.63 12.40
CV per capita (%) 2.135 0.459 1.103 1.130 0.291 0.529
8 Alternative Policy Simulations
Our model framework allows flexibility to explore alternative policy simulations. Recognizing that the actual implementation of a global ETS requires substantial international commitment, we consider two alternative policy scenarios to gain additional insights. First, we examine a partial ETS in which a major player in the world economy—the United States—does not participate, and the ETS operates among the remaining regions. Second, we analyze alternative net-zero pathways that vary in the stringency of global emission reduction commitments.
8.1 Partial ETS: US Non-participation
Recent political developments under the Trump administration have indicated that the United States is unlikely to participate in international cooperation aimed at mitigating global emissions in the near term. In light of this potential fragmentation of global climate cooperation, it is useful to numerically examine the economic and environmental consequences of a partial ETS. To this end, we simulate a scenario in which the United States refrains from any form of climate action—neither enforcing its emission cap nor participating in the ETS—while the remaining 11 regions continue to operate a cap-and-trade system.
Figure 6 compares the key outcomes under the baseline (“full ETS”) and the partial ETS scenarios. As intuitively expected, US non-participation leads to a substantial increase in global emissions and a corresponding rise in global temperature, reaching 1.83 degrees Celsius—about 0.12 degrees Celsius higher than in the baseline. Under the partial ETS scenario, the US MAC is equal to its SCC which rises to \$336 by 2100 due to higher global average temperatures.
Figure 6: Comparative analysis: effects of US non-participation in the global ETS
Table 3 shows welfare effects of the US non-participation by comparing the partial ETS to the full ETS, further illustrating the distributional effects of fragmented climate cooperation. The United States experiences a significant welfare gain, with CV per capita of \$565—approximately 1.3% of per capita consumption—reflecting its reduced abatement burden and the low level of MACs as shown in the figure above. On the other hand, the other regions (except for Russia) all experience significant welfare losses, ranging from -3.8% (MidEast) to -0.01% (OHI) of per capita consumption, as they bear the costs of both higher climate damages and continued emission constraints under the ETS. Russia is the only exception among the participating regions, experiencing welfare gains because higher global average temperatures are expected to benefit its economy, as noted earlier. Overall, this exercise underscores the critical importance of sustained and collective international commitment to climate mitigation, given the strong incentives for free-riding under noncooperative settings.
Table 3: Welfare effects of US non-participation in the global ETS
US EU Japan Russia Eurasia China
CV per capita (\$) 564.96 -54.00 -115.07 105.06 -129.73 -10.15
CV per capita (%) 1.309 -0.215 -0.404 1.226 -2.561 -0.103
India MidEast Africa LatAm OHI OthAs
CV per capita (\$) -28.94 -327.27 -52.56 -121.99 -3.23 -79.91
CV per capita (%) -1.867 -3.793 -3.409 -1.966 -0.009 -3.378
8.2 Net Zero Emission Scenarios
Along with the baseline emission cap scenario, we analyze simulation results for the noncooperative model with alternative emission cap paths, defined by the net zero emission targets in 2050, 2070, and 2090, where all regions achieve net zero emissions by the specified year. The top-left and top-right panels in Figure 7 display the emission permit prices and expected temperature increases under different emission cap scenarios. Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,621 per ton of carbon in 2049, and the temperature rise is restricted to 1.62 degrees Celsius by the end of this century. Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$616 and \$446 per ton of carbon in 2049, respectively. In these scenarios, the temperature rise by the end of the century is expected to reach 1.80 degrees Celsius and 1.99 degrees Celsius above the pre-industrial level, respectively. This result shows that the global target of restricting the temperature rise to 1.5 degrees Celsius is unattainable in a noncooperative world, even under the most restrictive net zero 2050 scenario, suggesting that stronger measures are needed to effectively regulate global emissions.
The bottom-left and bottom-right panels in Figure 7 show the regional MAC and SCC, taking the US as an example. The comparison of the regional SCC and MAC for all other regions are available in Appendix A.4.4, which show the same patterns as the US. Our results show that stricter emission caps lead to higher MACs. Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,493 per ton of carbon in 2048, compared to the peak of \$1,322 in 2082 under the net zero 2090 scenario. This is because more rigorous emission caps imposed on each region entail additional abatement efforts, resulting in a higher MAC. We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$225 per ton of carbon in 2050 in the net zero 2050 scenario, while it is \$270 in 2050 in the net zero 2090 scenario. This is because the permit price in the net zero 2050 scenario grows more quickly over time and even faster than the MAC.
@@ 9 Sensitivity Analysis
Lastly, recognizing that climate damages and emission abatement costs are key drivers of our model outcomes, including permit prices, MAC, and the SCC, we conduct sensitivity analyses on the parameters for climate damages ($\pi_{1,i}$, $\pi_{2,i}$) and emission abatement ($b_{1,i}$, $b_{2,i}$, $b_{3,i},$ $b_{4,i}$).
9.1 Climate Damage Parameters
@@
. While we incorporate climate damage projections from Burke et al., (2018) as our baseline model simulation, we additionally consider projections from Kahn et al., (2021) and Nordhaus, 2010a . Specifically, we calibrate the climate damage parameters to match the projections from Kahn et al., (2021) and directly adopt the parameters from Nordhaus, 2010a . Figure 8 demonstrates the key economic and climate outcomes under different climate damage assumptions. The baseline climate damage parameters from Burke et al., (2018) result in slightly lower emission permit prices, reaching \$2,105 by 2069, compared to \$2,322 under the damage parameters estimated from Kahn et al., (2021) and \$2,334 under the damage parameters from Nordhaus, 2010a . With the same emission cap constraints imposed, the global temperature outcomes remain unaffected despite variations in the damage parameters. For the US, the baseline climate damage estimation from Burke et al., (2018) leads to higher MAC and higher SCC, indicating that the baseline marginal climate damages are relatively higher than those projected by Kahn et al., (2021) and Nordhaus, 2010a . However, regional heterogeneity exists; for example, Russia experiences negative SCC under the baseline parameter values, while experiencing small but positive SCC under the parameter values estimated from Kahn et al., (2021) and Nordhaus, 2010a . See Appendix A.4.5 for a comparison across all 12 regions.
@@
9.2 Abatement Cost Parameters
For our sensitivity analysis of the emission abatement cost parameters, we consider the parameter values from Nordhaus, 2010a in addition to the baseline parameter values calibrated from Ueckerdt et al., (2019), both of which share the same functional form of abatement cost. As shown in Figure 9, the permit price rises to \$1,717 by 2069 under the abatement cost estimate of Nordhaus, 2010a , approximately 81% of the permit price projected under the baseline scenario. Despite the lower permit price, the global temperature increase remains similar to the baseline simulation, reaching 1.70 degrees Celsius by the end of the century, due to the emission cap constraints. The lower emission permit price under the abatement cost estimate of Nordhaus, 2010a reflects lower MAC, as illustrated with the US case in Figure 9, with similar patterns observed across most regions (see Appendix A.4.6). Lastly, the SCC is also lower under Nordhaus, 2010a parameters, with the SCC of the US reaching \$102 by 2100—just 33% of the baseline scenario.
@@ 10 Conclusion
In this work, we build a dynamic multi-region model of climate and the economy with a global emission cap-and-trade system. In our integrated assessment framework, regions participating in a global or partial ETS are allocated emission caps in line with the emission targets of the NDCs and net zero commitments, as established under the Paris Agreement and the Glasgow Pact. We solve for the market prices of emission permits under the dynamic Nash equilibrium in a noncooperative setting among the participating regions. The permit prices are endogenously determined by demand and supply of emission permits in the permits market, reflecting regional heterogeneity in future productivity growth, abatement technologies, climate damage, and population growth. For strictly positive emissions, we show both theoretically and numerically that the regional SCC is equal to the difference between the regional MAC and the market price of permits for regions participating in the ETS.
This work has several policy implications. First, our results indicate that the current global target of restricting the global temperature rise to 1.5 degrees Celsius above pre-industrial level by 2100 is unattainable under noncooperation, with the current emission commitments outlined in the Paris Agreement and the Glasgow Pact. Our findings suggest that more stringent emission reduction targets and global cooperation are needed to curb the trend of rising global temperature. Second, our baseline simulation shows that the current emission commitments lead to excess emission permit supply in the initial years, resulting in permit prices of zero. This finding suggests that effective implementation of the global ETS requires stricter emission caps so that the global supply of permits does not exceed the global demand of permits. Third, we demonstrate that an ETS is not a perfect substitute for a carbon tax; rather, the two instruments are complementary and can be jointly employed to enhance policy efficiency. Finally, our numerical analysis highlights the strong free-rider incentives in a partial ETS, illustrating why achieving sustained global cooperation in carbon mitigation remains a major challenge.
11 Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used generative AI (ChatGPT) in order to improve the manuscript’s readability during the revision process. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@ A.4.2 Model Comparison of ETS Implementation
Figure A.8 compares regional net emissions between two cases under noncooperation with the baseline emission caps: (i) with the ETS, (ii) without the ETS.
Figure A.8: Comparison of regional net emissions with and without the ETS.
@@
Figure A.9: Comparison of regional MAC with and without the ETS.
@@
Figure A.10: Comparison of regional SCC with and without ETS.
A.4.3 Alternative Policy Simulations: Partial ETS
Figure A.11 compares the regional MAC under noncooperation, comparing two cases: (i) with the full ETS, (ii) with a partial ETS.
Figure A.11: Comparison of regional MAC under full and partial ETS scenarios.
Figure A.12 compares the regional SCC under noncooperation, comparing two cases: (i) with the full ETS, (ii) with the partial ETS without participation of the United States.
Figure A.12: Comparison of regional SCC under the full and the partial ETS scenarios.
A.4.4 Alternative Policy Simulations: Net Zero Scenarios
@@
Figure A.14: Comparison of regional SCC under different emission caps.
A.4.5 Sensitivity Analysis over Climate Damage Parameters
In Figure A.15, we compare the MAC for the noncooperative model with the ETS under alternative climate damage parameters ($\pi_{1,i}$, $\pi_{2,i}$), based on projections from Kahn et al., (2021) and Nordhaus, 2010a alongside the baseline projection from Burke et al., (2018).
@@ A.4.5 Sensitivity Analysis over Climate Damage Parameters
A.4.6 Sensitivity Analysis over Abatement Cost Parameters
@@
A.4.7 Sensitivity over TFP Growth Rates